OpenStudy (anonymous):

How do you solve 8x^2+10x-7?

OpenStudy (anonymous):

We have \[8x^2+10x-7=0\] \[8x^2+14x-4x-7=0\] \[2x(4x+7)-1(4x+7)=0\] \[(2x-1)(4x+7)=0\] \[x=1/2\text{ or }x=-7/4\]